# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ B.1 Proofs of Score Network Approximation
Now we characterize the configuration of neural network $\widehat{f}(x,w,t)$. For boundedness, by Lemma A.10,
$\|\widehat{f}(x,w,t)\|\leq\|f_{*}\|_{L^{\infty}(\Omega_{R_{f}})}+\varepsilon\leq 2C_{X}^{\prime\prime}R_{f}^{6}=:M_{f}.$ (B.24)
Hence we can let $R_{f}=\mathcal{O}\left(\log^{\frac{1}{2}}\left(\frac{nK}{\varepsilon\delta}\right)\right)$ to ensure the lower bound of $R_{f}$ mentioned above and in Theorem A.8. For Lipschitzness, by (B.21),
$\|\widehat{f}(x,w,t)-\widehat{f}(x,\widetilde{w},t)\|$ $\lesssim\gamma_{w}(d_{x}+d_{y})\|w-\widetilde{w}\|_{\infty}$ (B.25)
@@
For the size of neural network, for each coordinate, by the construction in (B.18), the neural network $\widehat{g}_{i}$ consists of $N_{1}^{d_{x}+d_{y}}N_{2}$ parallel subnetworks, i.e., $g_{i}(u/N_{1},j/N_{2})\widehat{\Psi}_{u,j}(\cdot,\cdot,\cdot)$. By definition in (B.17), the subnetwork consists of $\mathcal{O}\left((d_{x}+d_{y})(d_{x}+d_{y}+\log\frac{R_{f}}{\varepsilon})\right)$ layers and the width is bounded by $\mathcal{O}(d_{x}+d_{y})$. Therefore, the whole neural network $\widehat{g}_{i}$ can be implemented by $\mathcal{O}\left((d_{x}+d_{y})(d_{x}+d_{y}+\log(R_{f}/\varepsilon))\right)$ layers with width $\mathcal{O}\left(N_{1}^{d_{x}+d_{y}}N_{2}(d_{x}+d_{y})\right)=\mathcal{O}\left(\frac{R_{f}^{3(d_{x}+d_{y})}}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right)$, and the number of parameter is bounded by $\mathcal{O}\left(\frac{R_{f}^{3(d_{x}+d_{y})}\log(R_{f}/\varepsilon)}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right)$. Combine these arguments together, we can claim that the size of neural network $\widehat{f}$ is
$L=\mathcal{O}\left((d_{x}+d_{y})(d_{x}+d_{y}+\log(R_{f}/\varepsilon))\right)=\mathcal{O}\left(\log\left(\frac{\log(nK/(\varepsilon\delta))}{\varepsilon}\right)\right),$ (B.26)
@@
$B=\mathcal{O}\left(N_{1}R_{f}+N_{2}T\right)=\mathcal{O}\left(\frac{R_{f}^{3}T}{\varepsilon}\right)=\mathcal{O}\left(\frac{T\log^{\frac{3}{2}}(nK/(\varepsilon\delta))}{\varepsilon}\right),$ (B.27)
@@
$\inf_{h\in\mathcal{H}}\|h-h_{*}\|_{L^{\infty}([0,1]^{D_{y}})}\leq\sqrt{d_{y}}\varepsilon,$ (B.28)
@@
$L_{h}=\mathcal{O}\left(\log(1/\varepsilon)\right),W_{h}=\mathcal{O}\left(\varepsilon^{-D_{y}}\log(1/\varepsilon)\right),$ (B.29)
@@
B.2 Proofs of Distribution Estimation
Theorem B.4 (Thm. 4.2).
@@ B.2 Proofs of Distribution Estimation
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D_{y}+2}}}+\sqrt{\Delta}.$ (B.30)
@@
Combine Theorem B.1 and Theorem A.6 and plug in the configuration of $\mathcal{F},\mathcal{H}$, we have with probability no less than $1-\delta$
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{0}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f}^{\mathbb{P}^{0}},\widehat{h}})]$ (B.31)
@@
By Lemma B.7,
$\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\sqrt{\mathbb{E}_{\mathbb{P}^{0}_{x|y}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f}^{\mathbb{P}^{0}},\widehat{h}})]}$ (B.32)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]$ $\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\nu^{-\frac{1}{2}}\log(nK/(\varepsilon\delta))\varepsilon+\sqrt{\Delta}$ (B.33)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]\lesssim\frac{\varepsilon_{0}}{\nu^{\frac{1}{2}}}+\sqrt{\Delta}+\frac{\log^{\frac{5(d_{x}+d_{y})+17}{4}}(\frac{nK}{\varepsilon_{0}\delta})\log^{\frac{3d_{x}+5}{2}}(\frac{1}{\varepsilon_{0}})}{(m\wedge(\nu n))^{\frac{1}{2}}\varepsilon_{0}^{\frac{d_{x}+d_{y}+7}{2}}}+\frac{\log^{2}(\frac{1}{\varepsilon_{0}})\log^{D_{y}+\frac{1}{2}}(\frac{nK}{\varepsilon_{0}\delta})}{(\nu nK)^{\frac{1}{2}}\varepsilon_{0}^{\frac{D_{y}}{2}}}.$ (B.34)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D_{y}+2}}}+\sqrt{\Delta}.$ (B.35)
Theorem B.5 (Thm. 4.3).
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]$ $\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D_{y}+2}}}.$ (B.36)
@@
Combine Theorem B.1 and Theorem A.8 and plug in the configuration of $\mathcal{F},\mathcal{H}$, we have with probability no less than $1-\delta$
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f}^{\mathbb{P}},\widehat{h}})]$ (B.37)
@@
By Lemma B.7,
$\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\sqrt{\mathbb{E}_{\mathbb{P}_{x|y}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f}^{\mathbb{P}},\widehat{h}})]}$ (B.38)
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]$ $\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\log(nK/(\varepsilon\delta))\varepsilon$ (B.39)
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]$ $\lesssim\varepsilon_{0}+\frac{\log^{\frac{5(d_{x}+d_{y})+17}{4}}(\frac{nK}{\varepsilon_{0}\delta})\log^{\frac{3d_{x}+5}{2}}(\frac{1}{\varepsilon_{0}})}{(m\wedge n)^{\frac{1}{2}}\varepsilon_{0}^{\frac{d_{x}+d_{y}+7}{2}}}$ (B.40)
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D_{y}+2}}}.$ (B.41)
B.3 Auxiliary Lemmas
Lemma B.6.
@@ B.3 Auxiliary Lemmas
According to (A.2),
$f_{*}^{\mathbb{P}}(x,w,t)=-\frac{x}{\sigma_{t}^{2}}+\frac{\alpha_{t}}{\sigma_{t}^{2}}\int x_{0}\frac{\phi_{t}(x|x_{0})p(x_{0};w)}{\int\phi_{t}(x|z)p(z;w)\mathrm{d}z}\mathrm{d}x_{0}.$ (B.42)
@@
$\frac{\partial}{\partial t}f_{*}^{\mathbb{P}}(x,w,t)=-\frac{2\alpha_{t}^{2}x}{\sigma_{t}^{2}}+\frac{\alpha_{t}}{\sigma_{t}^{2}}\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\left(x_{0},\frac{\partial}{\partial_{t}}\log\phi_{t}(x|x_{0})\right)-\frac{\alpha_{t}(1+\alpha_{t}^{2})}{\sigma_{t}^{4}}\mathbb{E}_{q_{t}(x_{0}|x,w)}[x_{0}].$ (B.43)
@@
$\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\left(x_{0},\frac{\partial}{\partial_{t}}\log\phi_{t}(x|x_{0})\right)$ $=-\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\left(x_{0},\frac{\partial}{\partial_{t}}\frac{\|x-\alpha_{t}x_{0}\|^{2}}{2\sigma_{t}^{2}}\right)$ (B.44)
@@
$\left\|\frac{\partial}{\partial t}f_{*}^{\mathbb{P}}(x,w,t)\right\|_{\infty}$ $\lesssim\frac{\alpha_{t}^{2}R_{f}}{\sigma_{t}^{2}}+\mathbb{E}_{q_{t}(x_{0}|x,w)}\Big{\|}\frac{x-\alpha_{t}x_{0}}{\sigma_{t}^{2}}\Big{\|}^{3}+\frac{\alpha_{t}(1+\alpha_{t}^{2})}{\sigma_{t}^{4}}\mathbb{E}_{q_{t}(x_{0}|x,w)}[\|x_{0}\|_{\infty}]$ (B.45)
@@
$\mathbb{E}_{q_{t}(x_{0}|x,w)}\Big{\|}\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big{\|}^{3}$ $\preceq\frac{1}{\sigma_{t}^{3}}\int\big{\|}\frac{\alpha_{t}x_{0}-x}{\sigma_{t}}\big{\|}^{3}\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}$ (B.46)
@@
$\left\|\frac{\partial}{\partial t}f_{*}^{\mathbb{P}}(x,w,t)\right\|_{\infty}$ $\lesssim\frac{R_{f}^{2}}{\sigma_{t}^{2}}+\frac{R_{f}^{3}+C_{0}^{3}}{\sigma_{t}^{6}}+\frac{R_{f}+C_{0}}{\sigma_{t}^{3}}$ $\lesssim\frac{R_{f}^{3}}{T_{0}^{3}}.$ (B.48)
Lemma B.7.
@@
$\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\sqrt{\mathbb{E}_{\mathbb{P}^{0}_{x|y}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f},\widehat{h}})]}.$ (B.49)
@@
$\displaystyle d\widetilde{x}_{t}=(\widetilde{x}_{t}+2\nabla\log p_{T-t}(\widetilde{x}_{t}|y))\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},\ \widetilde{x}_{0}\sim\mathcal{N}(0,I),0\leq t\leq T-T_{0},$ (B.50)
$\displaystyle d\bar{x}_{t}=(\bar{x}_{t}+2\nabla\log p_{T-t}(\widetilde{x}_{t}|y))\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},\ \bar{x}_{0}\sim p_{T},0\leq t\leq T-T_{0}.$ (B.51)
Denote the distribution of $\widetilde{x}_{t}$ as ~$\widetilde{\mathbb{P}}_{T-t}$. And note that $\bar{x}_{t}\sim p_{T-t}$ by classic reverse-time SDE results [Anderson, 1982]. Then by Fu et al. [2024, Lemma D.5],
$\mathrm{TV}(\mathbb{P}_{T_{0}},\mathbb{P}_{0})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0}).$ (B.52)

### v2

@@ C.1 Proofs of Score Network Approximation
Now we characterize the configuration of neural network $\widehat{f}(x,w,t)$. For boundedness, by Lemma B.10,
$\|\widehat{f}(x,w,t)\|\leq\|f_{*}\|_{L^{\infty}(\Omega_{R_{f}})}+\varepsilon\leq 2C_{X}^{\prime\prime}R_{f}^{6}=:M_{f}.$ (C.24)
Hence we can let $R_{f}=\mathcal{O}\left(\log^{\frac{1}{2}}\left(\frac{nK}{\varepsilon\delta}\right)\right)$ to ensure the lower bound of $R_{f}$ mentioned above and in Theorem B.8. For Lipschitzness, by (C.21),
$\|\widehat{f}(x,w,t)-\widehat{f}(x,\widetilde{w},t)\|$ $\lesssim\gamma_{w}(d_{x}+d_{y})\|w-\widetilde{w}\|_{\infty}$ (C.25)
@@
For the size of neural network, for each coordinate, by the construction in (C.18), the neural network $\widehat{g}_{i}$ consists of $N_{1}^{d_{x}+d_{y}}N_{2}$ parallel subnetworks, i.e., $g_{i}(u/N_{1},j/N_{2})\widehat{\Psi}_{u,j}(\cdot,\cdot,\cdot)$. By definition in (C.17), the subnetwork consists of $\mathcal{O}\left((d_{x}+d_{y})(d_{x}+d_{y}+\log\frac{R_{f}}{\varepsilon})\right)$ layers and the width is bounded by $\mathcal{O}(d_{x}+d_{y})$. Therefore, the whole neural network $\widehat{g}_{i}$ can be implemented by $\mathcal{O}\left((d_{x}+d_{y})(d_{x}+d_{y}+\log(R_{f}/\varepsilon))\right)$ layers with width $\mathcal{O}\left(N_{1}^{d_{x}+d_{y}}N_{2}(d_{x}+d_{y})\right)=\mathcal{O}\left(\frac{R_{f}^{3(d_{x}+d_{y})}}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right)$, and the number of parameter is bounded by $\mathcal{O}\left(\frac{R_{f}^{3(d_{x}+d_{y})}\log(R_{f}/\varepsilon)}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right)$. Combine these arguments together, we can claim that the size of neural network $\widehat{f}$ is
$L=\mathcal{O}\left((d_{x}+d_{y})(d_{x}+d_{y}+\log(R_{f}/\varepsilon))\right)=\mathcal{O}\left(\log\left(\frac{\log(nK/(\varepsilon\delta))}{\varepsilon}\right)\right),$ (C.26)
@@
$B=\mathcal{O}\left(N_{1}R_{f}+N_{2}T\right)=\mathcal{O}\left(\frac{R_{f}^{3}T}{\varepsilon}\right)=\mathcal{O}\left(\frac{T\log^{\frac{3}{2}}(nK/(\varepsilon\delta))}{\varepsilon}\right),$ (C.27)
@@
$\inf_{h\in\mathcal{H}}\|h-h_{*}\|_{L^{\infty}([0,1]^{D_{y}})}\leq\sqrt{d_{y}}\varepsilon,$ (C.28)
@@
$L_{h}=\mathcal{O}\left(\log(1/\varepsilon)\right),W_{h}=\mathcal{O}\left(\varepsilon^{-D_{y}}\log(1/\varepsilon)\right),$ (C.29)
@@
C.2 Proofs of Distribution Estimation
Theorem C.4 (Thm. 4.2).
@@ C.2 Proofs of Distribution Estimation
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D_{y}+2}}}+\sqrt{\Delta}.$ (C.30)
@@
Combine Theorem C.1 and Theorem B.6 and plug in the configuration of $\mathcal{F},\mathcal{H}$, we have with probability no less than $1-\delta$
[…]
@@
By Lemma C.7,
$\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\sqrt{\mathbb{E}_{\mathbb{P}^{0}_{x|y}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f}^{\mathbb{P}^{0}},\widehat{h}})]}$ (C.32)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]$ $\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\nu^{-\frac{1}{2}}\log(nK/(\varepsilon\delta))\varepsilon+\sqrt{\Delta}$ (C.33)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]\lesssim\frac{\varepsilon_{0}}{\nu^{\frac{1}{2}}}+\sqrt{\Delta}+\frac{\log^{\frac{5(d_{x}+d_{y})+17}{4}}(\frac{nK}{\varepsilon_{0}\delta})\log^{\frac{3d_{x}+5}{2}}(\frac{1}{\varepsilon_{0}})}{(m\wedge(\nu n))^{\frac{1}{2}}\varepsilon_{0}^{\frac{d_{x}+d_{y}+7}{2}}}+\frac{\log^{2}(\frac{1}{\varepsilon_{0}})\log^{D_{y}+\frac{1}{2}}(\frac{nK}{\varepsilon_{0}\delta})}{(\nu nK)^{\frac{1}{2}}\varepsilon_{0}^{\frac{D_{y}}{2}}}.$ (C.34)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}^{0}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D_{y}+2}}}+\sqrt{\Delta}.$ (C.35)
Theorem C.5 (Thm. 4.3).
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]$ $\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D_{y}+2}}}.$ (C.36)
@@
Combine Theorem C.1 and Theorem B.8 and plug in the configuration of $\mathcal{F},\mathcal{H}$, we have with probability no less than $1-\delta$
[…]
@@
By Lemma C.7,
$\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\sqrt{\mathbb{E}_{\mathbb{P}_{x|y}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f}^{\mathbb{P}},\widehat{h}})]}$ (C.38)
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]$ $\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\log(nK/(\varepsilon\delta))\varepsilon$ (C.39)
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]$ $\lesssim\varepsilon_{0}+\frac{\log^{\frac{5(d_{x}+d_{y})+17}{4}}(\frac{nK}{\varepsilon_{0}\delta})\log^{\frac{3d_{x}+5}{2}}(\frac{1}{\varepsilon_{0}})}{(m\wedge n)^{\frac{1}{2}}\varepsilon_{0}^{\frac{d_{x}+d_{y}+7}{2}}}$ (C.40)
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{y\sim\mathbb{P}_{y}}[\mathrm{TV}(\widehat{\mathbb{P}}_{x|y},\mathbb{P}_{x|y})]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{d_{x}+d_{y}+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D_{y}+2}}}.$ (C.41)
C.3 Auxiliary Lemmas
Lemma C.6.
@@ C.3 Auxiliary Lemmas
According to (B.2),
$f_{*}^{\mathbb{P}}(x,w,t)=-\frac{x}{\sigma_{t}^{2}}+\frac{\alpha_{t}}{\sigma_{t}^{2}}\int x_{0}\frac{\phi_{t}(x|x_{0})p(x_{0};w)}{\int\phi_{t}(x|z)p(z;w)\mathrm{d}z}\mathrm{d}x_{0}.$ (C.42)
@@
$\frac{\partial}{\partial t}f_{*}^{\mathbb{P}}(x,w,t)=-\frac{2\alpha_{t}^{2}x}{\sigma_{t}^{2}}+\frac{\alpha_{t}}{\sigma_{t}^{2}}\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\left(x_{0},\frac{\partial}{\partial_{t}}\log\phi_{t}(x|x_{0})\right)-\frac{\alpha_{t}(1+\alpha_{t}^{2})}{\sigma_{t}^{4}}\mathbb{E}_{q_{t}(x_{0}|x,w)}[x_{0}].$ (C.43)
@@
$\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\left(x_{0},\frac{\partial}{\partial_{t}}\log\phi_{t}(x|x_{0})\right)$ $=-\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\left(x_{0},\frac{\partial}{\partial_{t}}\frac{\|x-\alpha_{t}x_{0}\|^{2}}{2\sigma_{t}^{2}}\right)$ (C.44)
@@
$\left\|\frac{\partial}{\partial t}f_{*}^{\mathbb{P}}(x,w,t)\right\|_{\infty}$ $\lesssim\frac{\alpha_{t}^{2}R_{f}}{\sigma_{t}^{2}}+\mathbb{E}_{q_{t}(x_{0}|x,w)}\Big\|\frac{x-\alpha_{t}x_{0}}{\sigma_{t}^{2}}\Big\|^{3}+\frac{\alpha_{t}(1+\alpha_{t}^{2})}{\sigma_{t}^{4}}\mathbb{E}_{q_{t}(x_{0}|x,w)}[\|x_{0}\|_{\infty}]$ (C.45)
@@
$\mathbb{E}_{q_{t}(x_{0}|x,w)}\Big\|\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big\|^{3}$ $\preceq\frac{1}{\sigma_{t}^{3}}\int\big\|\frac{\alpha_{t}x_{0}-x}{\sigma_{t}}\big\|^{3}\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}$ (C.46)
@@
$\int_{\|\frac{\alpha_{t}x_{0}-x}{\sigma_{t}}\|\leq R/2}p(x_{0};w)\mathrm{d}x_{0}\geq\int_{\|x_{0}\|\leq C_{0}}p(x_{0};w)\mathrm{d}x_{0}\geq 1-2\exp(-C_{1}^{\prime}C_{0}^{2})\geq\frac{1}{2},$ (C.47)
@@
$\left\|\frac{\partial}{\partial t}f_{*}^{\mathbb{P}}(x,w,t)\right\|_{\infty}$ $\lesssim\frac{R_{f}^{2}}{\sigma_{t}^{2}}+\frac{R_{f}^{3}+C_{0}^{3}}{\sigma_{t}^{6}}+\frac{R_{f}+C_{0}}{\sigma_{t}^{3}}$ $\lesssim\frac{R_{f}^{3}}{T_{0}^{3}}.$ (C.48)
Lemma C.7.
@@
$\mathrm{TV}(\widehat{\mathbb{P}}^{0}_{x|y},\mathbb{P}^{0}_{x|y})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0})+e^{-T}+\sqrt{\mathbb{E}_{\mathbb{P}^{0}_{x|y}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f},\widehat{h}})]}$. (C.49)
@@
$\displaystyle d\widetilde{x}_{t}=(\widetilde{x}_{t}+2\nabla\log p_{T-t}(\widetilde{x}_{t}|y))\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},\ \widetilde{x}_{0}\sim\mathcal{N}(0,I),0\leq t\leq T-T_{0},$ (C.50)
$\displaystyle d\bar{x}_{t}=(\bar{x}_{t}+2\nabla\log p_{T-t}(\widetilde{x}_{t}|y))\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},\ \bar{x}_{0}\sim p_{T},0\leq t\leq T-T_{0}.$ (C.51)
Denote the distribution of $\widetilde{x}_{t}$ as $\widetilde{\mathbb{P}}_{T-t}$. And note that $\bar{x}_{t}\sim p_{T-t}$ by classic reverse-time SDE results [Anderson, 1982]. Then by Fu et al. [2024, Lemma D.5],
$\mathrm{TV}(\mathbb{P}_{T_{0}},\mathbb{P}_{0})\lesssim\sqrt{T_{0}}\log^{\frac{d_{x}+1}{2}}(1/T_{0}).$ (C.52)